A cyclist travels 26.7 km in 54.0 min 16.0 s. What is his average velocity in m/s?
Answer #1

here,

the time interval , t = 54 min 16 s

t = 54 * 60 s + 16 s = 3256 s

the distance traveled , s = 26.7 km

s = 26700 m

the average velocity , v = s /t

v = 26700 /(3256) m/s

v = 8.2 m/s